OpenStudy (anonymous):

Let S be the parabolid x^2+y^2+z = R^2 , 0

OpenStudy (anonymous):

I'm gonna use the formula : \[\int\limits F (x, y, f(x,y)) * (-f_x - f_y +k ) dx dy \]

OpenStudy (anonymous):

so I have \[(−fx−fy+k)dxdy = (-2 x i - 2yj +k)dxdy\]

OpenStudy (anonymous):

so what would \[F(x,y,f(x,y))\] be ?

OpenStudy (anonymous):

\[F(x i + yj + (-x-y+R)k) ???\]

OpenStudy (anonymous):

\[\nabla\]

OpenStudy (jhannybean):

Thank you!

OpenStudy (anonymous):

so I take the gradient F?

OpenStudy (jhannybean):

\[\vec F = \nabla f\] yes.

TheSmartOne (thesmartone):

\(\ \nabla\) @Jhannybean :P

TheSmartOne (thesmartone):

nevermind...

OpenStudy (anonymous):

gradient is i + j + 2zk

OpenStudy (anonymous):

so i sub in \[F⃗ =∇f\] on my formula?

OpenStudy (jhannybean):

You are like 3 sections ahead of me!! Haha, I'm working on Stokes' and Divergent Theorem atm :(

OpenStudy (anonymous):

meet too

OpenStudy (anonymous):

i suppose to use the general formula to find the flux for this problem

OpenStudy (jhannybean):

Isn't flux : \[\int\int_S \vec F \cdot dS = \int\int_S = \vec F \cdot \vec n dS\]

OpenStudy (jhannybean):

\[\vec n = \frac{|\nabla f |}{\sqrt{| \nabla f|}}\]

OpenStudy (anonymous):

yeah but why would i need the unit normal?

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

et S be the parabolid x^2+y^2+z +R^2 , 0<z<R^2, oriented upward and let F = xi + yj +z^2k. find the flux of the vector field F through the surface S.

OpenStudy (dan815):

x^2+y^2+z =R^2 ?

OpenStudy (anonymous):

yeah = R^2 sorry

OpenStudy (anonymous):

I edit the question

OpenStudy (dan815):

ok

OpenStudy (dan815):

do div F dv

OpenStudy (dan815):



OpenStudy (dan815):

uh

OpenStudy (dan815):

wrong surface

OpenStudy (dan815):

-z=x^2+y^2-r^2

OpenStudy (dan815):